# Quantum-gas microscopy of the Bose-glass phase

### v1

@@ Abstract
Figure 3: Talbot Interferometry. Coherence length, $\xi$, in units of the lattice spacing, $\,a_{\rm{\ell}}$, for increasing disorder strength ($V_{x,y}=9\,E_{\rm{r}}$, $U/J=11$). The coherence length is extracted from the fitted decay time of the Talbot interferometry measurements (insets) by comparing to a numerical calculation (Methods). Error bars represent the $68\%$ confidence intervals of the fitted decay times. Each data point of the curves in the insets results from the average of five images, each containing $\sim 200$ atoms. The error bars are the standard error. Black lines are fits with a damped sine, the decay time of which is used to extract $\xi$ (Methods).
@@
Figure 4: Ergodicity and adiabaticity. a Sketch of the phase diagram (not to scale) of the disordered Bose-Hubbard model interpreted from the datasets in Fig. 2. The blue circles show the points and the arrows the trajectories used for the data sets shown in b and c. b Coherence length, $\xi$, measured using Talbot interferometry for the points and trajectories shown in a (details see text). Each data point results from fits of four repetitions with the same parameters. c EA parameter, $q$, for three points and trajectories as shown in a. Each measurement is an average over four disorder patterns with ten images each. d Cloud width after one Talbot time, $\tau_{\rm T}$, after changing from the Mott-insulator regime to superfluid regime for increasing disorder strength (see inset where the dashed arrow indicates the trajectory for a data point at a particular disorder strength).
@@
We employ this method to quantitatively probe how the coherence length $\xi$ of an initially superfluid system is affected by the addition of disorder. To measure the cloud width, we calculate the average distance of each atom from the center, after a thermalization time of $250\,\mathrm{ms}$ at $V_{x,y}=6\,E_{\rm{r}}$. The disorder potential is switched off simultaneously with the lattice and remains off during thermalization and imaging. The resulting curve shows a damped oscillation as a function of $t$ (insets in Fig. 3), where the presence of $n$ half-periods of oscillation indicates phase coherence over $n$ lattice sites [39]. Without disorder, the oscillation shows a weak decay with all peaks pronounced, indicating long-range phase coherence. When adding disorder, we observe that the oscillations decay more rapidly. At the maximum experimentally accessible disorder strength, only the first two peaks are distinguishable from zero (insets in Fig. 3). To quantify the coherence length, we compare the fitted decay time of the oscillations to a numerical calculation, using a density matrix $\rho(i,j)\sim\exp{(-\lvert i-j\rvert/\xi)}$ with exponentially decaying coherence (Methods). While for $\Delta=0$ we observe a coherence length $\xi=15(3)\,a_{\rm{\ell}}$, the coherence length decreases to $\xi=1.4(3)\,a_{\rm{\ell}}$ for our maximum available disorder strength (Fig. 3). This observation is consistent with the characteristics of a Bose glass, which is expected to exhibit coherence locally in small regions but not across the entire system [15, 45]. Our analysis shows the advantage of the Talbot interferometry measurements to give a quantitative measurement of the coherence length, in contrast to the time-of-flight visibility (Fig. 2). In the Mott-insulating phase, the oscillations from the Talbot interferometry measurements are significantly more damped (Supplementary Materials) compared to those shown in Fig. 3, indicating a coherence length smaller than a lattice site.
Ergodicity, adiabaticity, and coherence
@@
In a finite-size system, a phase transition can be crossed without significant heating if both phases are ergodic. We explore the interesting question of how the Mott-insulator-to-superfluid transition is affected by the presence of the Bose glass. Signatures of non-ergodicity in the Bose-glass-to-superfluid transition were recently observed using a quasicrystalline potential [31]. To probe adiabaticity, we cross from a superfluid or Mott insulator to a parameter region where we have identified the presence of the Bose glass (in Figs. 2b and c) by changing the lattice depth while keeping the same disorder strength and then attempt to restore the original state by reversing the procedure.
@@
Initially, we perform Talbot interferometry for a set of points in the superfluid phase ($U/J=5.5$), for three different disorder strengths (points $a$,$b$ and $c$ in Fig. 4a), revealing similar long coherence lengths (Fig. 4b). We also verify that transitioning to the Mott insulator and back to the superfluid at $\Delta=0$ (trajectory $a_{1}$ in Fig. 4a) restores a comparable coherence length (Fig. 4b). This was done by increasing the lattice depth to reach $U/J=25.5$ within $200\,\mathrm{ms}$ and vice versa. Next, we investigate how entering the Bose glass affects the system’s ability to restore phase coherence. Starting with a superfluid with weak disorder and increasing the lattice depths to reach $U/J=10.8$ in $200\,\mathrm{ms}$ and vice versa (trajectory $b_{1}$ in Fig. 4a), we observe a significantly reduced coherence length. We conduct two additional measurements, starting from a superfluid with stronger initial disorder to reach $U/J=14.6$ and $U/J=25.5$ (trajectories $c_{1}$ and $c_{2}$ in Fig. 4a).
@@
These measurements show a further reduction in coherence length (Fig. 4b), consistent with a larger proportion of the cloud entering the Bose glass compared to $b_{1}$, indicating the inability of the Bose-glass component to adiabatically transition back into the superfluid. The observed loss of coherence cannot be attributed to intensity noise on the lattice lasers as we were able to restore coherence when transitioning from the Mott insulator to the superfluid (trajectory $a_{1}$). We have verified that the coherence does not change when extending the duration over which we change the lattice depth from $200\,\mathrm{ms}$ to $500\,\mathrm{ms}$, which is the longest accessible timescale in our experiment.
@@
Similarly, we explore the transition from the Bose-glass phase to the Mott insulator. We first measure the Edwards-Anderson parameter in a Mott insulator in the presence of weak disorder ($U/J=110$, $\Delta/U_{c}=1.1$, point $d$ in Fig. 4a), and find $q\approx 0$ in the center of the cloud (Fig. 4c). When we decrease the lattice depth to reach $U/J=55$ and return to $U/J=110$ (trajectory $d_{1}$), we observe an increase of $q$ in the outer region of the cloud, which we attribute to the Bose glass having formed here due to the lower chemical potential caused by the harmonic confinement. We then perform a similar measurement, changing instead to $U/J=25.5$ and back (trajectory $d_{2}$), to cross fully into the Bose-glass regime. When attempting to transition back to the original state, we observe an increase of $q$ over the entire cloud (Fig. 4c), again indicating non-adiabaticity.
@@
Lastly, we investigate the effect of crossing the Bose glass at increasing disorder strengths, starting in the Mott insulator and finishing in the superfluid (see inset in Fig. 4d). Specifically, we prepare the system at $U/J=110$ and transition to $U/J=5.5$ within 200 ms, before using Talbot interferometry to measure the cloud width after $\tau_{\rm T}$, at the first minimum of the oscillation. We observe an increase of the width for increasing disorder strengths (Fig. 4d), indicating a loss of coherence. We attribute this to excitations in the system due to the expected non-ergodicity [31] when crossing an increasingly large Bose-glass region in the phase diagram.
Conclusion
@@
In summary, we identified the Bose-glass phase experimentally in both weakly and strongly interacting regimes in a two-dimensional square lattice with controllable and reproducible disorder potentials. Importantly, using Talbot interferometry across the superfluid-to-Bose-glass transition, we observed a change from long to short-range coherence. Additionally, we found that transitioning in and out of the Bose glass is non-adiabatic on the accessible experimental timescales. In future studies it may be possible to cross transitions adiabatically on longer timescales, by reducing trap losses and heating, via ultra-low vacuum pressure [46], and optical lattices with reduced intensity noise. This would enable the exploration of ergodic and non-ergodic phases in a single model, for both weak and strong interactions, through scaling of the adiabatic timescales for traversing transitions or restoring system properties.
@@
Our low-energy studies, combined with Talbot interferometry provide a starting point to further explore disordered systems in and out of equilibrium, and particularly to probe the timescales to establish phase coherence in these systems. This is of relevance to understanding disordered and glass-like quantum states which have been observed in, e.g., cavity polaritons [47], quantum magnets [48] and overdoped cuprates [49]. Future studies with ultracold atoms could focus on the influence of disorder in different lattice geometries [50, 51] and lower dimensions [52, 53], or further explore the stability of many-body localization [29] and the onset of quantum avalanches [54]. Furthermore, controlled disorder is important in studying frustration and stability of the spin-glass phase in the Fermi-Hubbard model [55, 56]. By extending the system sizes is also possible to study the similarities and differences between disordered and quasicrystalline systems, e.g., by realization of the 2D extension of the Aubry-Andre model [57, 58], which could feature mobility edges, critical states and give further insights into the nature of many-body localization and the Bose glass [31].
Methods
@@
We use the same quantum-gas microscope setup as detailed in our previous work [59]. Initially, we prepare a small Bose-Einstein condensate of $\sim 200$ atoms in a single antinode of an optical lattice (wavelength $\lambda=1064\,\mathrm{nm}$) in the vertical direction, with a depth of $V_{z}=25\,E_{\rm{r}}$. Here, $\,E_{\rm{r}}=h^{2}/2m\lambda^{2}$ is the recoil energy and $m$ is the mass of a 87Rb atom. To prepare a disordered system in the weakly interacting regime, we turn on the horizontal lattices, $V_{x}$ and $V_{y}$, to the desired depth, $V_{f}\leq 10\,E_{\rm{r}}$, using an s-shaped ramp within 200 ms and then turn on the disorder potential within 200 ms again using an s-shaped ramp. To prepare a disordered system in the strongly interacting regime, we first turn on the horizontal lattices to reach $V_{x,y}=10\,E_{\rm{r}}$, before adding the disorder while simultaneously increasing $V_{x}$ and $V_{y}$ to their final values, $V_{f}>10\,E_{\rm{r}}$ in a further $200\,\mathrm{ms}$. We detect the atoms by collecting their fluorescence light with a high-NA microscope objective and use the Lucy-Richardson deconvolution method [60] to infer the site occupations.
@@
We create the repulsive disorder potentials using a laser ($\lambda=666\,\mathrm{nm}$) incident on a digital micromirror device (DMD), similar to our previous work [59]. The light from the DMD is overlapped with the optical path used for fluorescence imaging of the atoms and projected through the microscope onto the atoms in the 2D plane. The magnification is such that approximately $18\times 18$ pixels on the DMD correspond to one lattice site. We create a potential at each lattice site (pictured in Fig. 1c), with random amplitude $\Delta_{i}\in[0,\Delta]$, that follows a uniform box distribution. The disorder pattern covers a region with a diameter of 25 lattice sites.
A motorized flip mirror can direct the light pattern from the DMD onto a monitor camera. We employ an iterative intensity feedback algorithm to shape the Gaussian beam incident on the DMD into the target disorder pattern. The light pattern is blurred by the microscope objective when it is projected onto the atoms. To compensate for this, we convolve the monitor camera images with the point-spread function of the objective to effectively optimize the potential as seen by the atoms. Once the DMD pattern has been calculated, it is saved and displayed continuously. The disorder strength is controlled solely by changing the laser power. As detailed in our previous work [59], we measure the phase drift of the optical lattice between experimental realizations by analyzing the position of single atoms in the fluorescence image. This information is then used to shift the position of the DMD pattern for the next measurement, ensuring that the disorder potential remains aligned with the optical lattice. The estimated phase drift from one experimental realization to the next is $<0.05\,a_{\rm{\ell}}$.
@@
Experimentally, the thermalization in two dimensions after the one-dimensional Talbot evolution averages over the different chain lengths in our system. We numerically checked that explicitly adding the energy of the one-dimensional chains of different length results in the same behavior as Eq. 6.
Acknowledgments
Supplementary Materials
@@ 4 Additional data: Talbot interferometry
5 Additional data: Entering the Boseglass phase

### v2

@@ 5 Talbot interferometry
Figure 3: Talbot interferometry. Coherence length, $\xi$, in units of the lattice spacing, $a_{\rm{\ell}}$, for increasing disorder strength at constant lattice depth ($V_{x,y}=9\,E_{\rm{r}}$, $U/J=11$). The coherence length is extracted from the fitted decay time of the Talbot interferometry measurements (insets) by comparing to a numerical calculation (Appendix). Error bars represent the $68\%$ confidence intervals of the fitted decay times. Each data point of the curves in the insets results from the average of five images, each containing $\sim 200$ atoms. The error bars are the standard error. Black lines are fits with a damped sine, the decay time of which is used to extract $\xi$ (Appendix).
@@
Figure 4: Ergodicity and adiabaticity. a Approximate phase diagram of the disordered Bose-Hubbard model interpreted from the datasets in Fig. 2. The blue shaded area highlights the region where the entire cloud is in the superfluid phase and the red shaded region shows where the center of the cloud is Mott-insulating. The blue circles show the points and the arrows the trajectories used for the data sets shown in b and c. b Coherence length, $\xi$, measured using Talbot interferometry for the points and trajectories shown in a (details see text). Each data point results from fits of four repetitions with the same parameters. c EA parameter, $q$, for three points and trajectories as shown in a. Each measurement is an average over four disorder patterns with ten images each. d Cloud width after one Talbot time, $\tau_{\rm T}$, after a ramp from the Mott-insulator regime to superfluid regime for increasing disorder strength (see inset where the dashed arrow indicates the trajectory for a data point at a particular disorder strength).
@@
We employ this method to quantitatively probe how the coherence length, $\xi$, of an initially superfluid system is affected by the addition of disorder. To measure the cloud width, we calculate the average distance of each atom from the center, after a thermalization time of $250\,\mathrm{ms}$ at $V_{x,y}=6\,E_{\rm{r}}$. The disorder potential is switched off simultaneously with the lattice and remains off during thermalization and imaging. The resulting curve shows a damped oscillation as a function of $t$, (insets in Fig. 3), where the presence of $n$ half-periods of oscillation indicates phase coherence over $n$ lattice sites [54]. For $\Delta=0$, the oscillation shows a weak decay with all peaks pronounced, indicating long-range phase coherence. When adding disorder, we observe that the oscillations decay more rapidly. At the maximum experimentally accessible disorder strength, only the first two peaks are distinguishable from zero. To quantify the coherence length, we fit the data with an exponentially damped oscillation (solid lines in the insets of Fig. 3). We compare the fitted decay times to a numerical calculation, which use a density matrix, $\rho(i,j)\sim\exp{(-\lvert i-j\rvert/\xi)}$, with exponentially decaying coherence, see Appendix. While for $\Delta=0$ we observe a coherence length $\xi=15(3)\,a_{\rm{\ell}}$, the coherence length decreases to $\xi=1.4(3)\,a_{\rm{\ell}}$ for our maximum available disorder strength, see Fig. 3. This observation is consistent with the characteristics of a Bose glass, which is expected to exhibit coherence locally in small regions but not across the entire system [24, 27]. Our analysis shows the advantage of the Talbot interferometry measurements to give a quantitative measurement of the coherence length, in contrast to the time-of-flight visibility (Fig. 2). In the Mott-insulating phase, the oscillations from the Talbot interferometry measurements are significantly more damped (Supplemental Material) compared to those shown in Fig. 3, indicating a coherence length smaller than a lattice site.
6 Ergodicity, adiabaticity, and coherence
@@
In a finite-size system, a phase transition can be crossed without significant heating if both phases are ergodic. Signatures of non-ergodicity in the Bose-glass-to-superfluid transition were recently observed using a quasicrystalline potential [68]. Here, we explore in a square lattice how the Mott-insulator-to-superfluid transition is affected by the presence of disorder. Using the data shown in Figs. 2b and c, we draw an approximate phase diagram for our system (Fig. 4a) to serve as a guide for the subsequent measurements.
Initially, we perform Talbot interferometry for a set of points in the superfluid phase ($U/J=5.5$), for three different disorder strengths. We find that the addition of the disorder potential does not result in significant heating by holding a disordered superfluid for $400$ ms (points $b$ and $c$ in Fig. 4a) and observe long-range coherence comparable to the case without disorder (point $a$). Importantly, this relates back to our earlier measurements, where the reduced visibility at large disorder strengths (Fig. 2b), as well as the quantitative observation of the reduction of coherence length in the Talbot measurements (Fig. 3), can then be interpreted as signatures of the emerging Bose glass.
@@
To probe adiabaticity, we first verify that transitioning to the Mott insulator and back to the superfluid at $\Delta=0$ (trajectory $a_{1}$ in Fig. 4a) restores a comparable coherence length, shown in Fig. 4b. This was done by increasing $V_{x}$ and $V_{y}$ to reach $U/J=25.5$ in $200\,\mathrm{ms}$ and vice versa. Next, we investigate how entering the region where we identified the Bose glass affects the system’s ability to restore phase coherence. Starting with a superfluid with weak disorder and increasing $V_{x}$ and $V_{y}$ to reach $U/J=10.8$ in $200\,\mathrm{ms}$ and vice versa (trajectory $b_{1}$ in Fig. 4a) we observe a significantly reduced coherence length. We conduct two additional measurements, starting from a superfluid with stronger initial disorder to reach $U/J=14.6$ and $U/J=25.5$ (trajectories $c_{1}$ and $c_{2}$ in Fig. 4a). These measurements show a further reduction in coherence length, shown in Fig. 4b, consistent with a larger proportion of the cloud entering the Bose glass compared to $b_{1}$, indicating the inability of the Bose-glass component to adiabatically transition back into the superfluid.
@@
The observed loss of coherence cannot be attributed to intensity noise on the lattice lasers as we were able to restore the initial coherence when transitioning from the Mott insulator to the superfluid (trajectory $a_{1}$). We have also verified that the coherence does not change when extending the duration over which we change $V_{x}$ and $V_{y}$ for the presented trajectories from $200\,\mathrm{ms}$ to $500\,\mathrm{ms}$, which is the longest accessible timescale in our experiment.
@@
Next, we explore the transition from the identified Bose-glass region to the Mott insulator. We first measure the Edwards-Anderson parameter in a Mott insulator in the presence of weak disorder ($U/J=110$, $\Delta/U_{c}=1.1$, point $d$ in Fig. 4a), and find $q\approx 0$ in the center of the cloud, as shown in Fig. 4c. When we decrease the lattice depth to reach $U/J=55$ and return to $U/J=110$ (trajectory $d_{1}$), we observe an increase of $q$ in the outer region of the cloud, which we attribute to localization occurring here first due to the lower chemical potential caused by the harmonic confinement. We then perform a similar measurement, changing instead to $U/J=25.5$ and back (trajectory $d_{2}$), to cross fully into the Bose-glass regime. When attempting to transition back to the original state, we observe an increase of $q$ over the entire cloud, see Fig. 4c, again indicating non-adiabaticity.
@@
Lastly, we investigate the effect of crossing the Bose glass at increasing disorder strengths, starting in the Mott insulator and finishing in the superfluid, see inset in Fig. 4d. Specifically, we prepare the system at $U/J=110$ and transition to $U/J=5.5$ within 200 ms, before using Talbot interferometry to measure the cloud width after $\tau_{\rm T}$, at the first minimum of the oscillation. We observe an increase in cloud width for increasing disorder strengths, shown in Fig. 4d, indicating a loss of coherence. We attribute this to excitations in the system due to the expected non-ergodicity [68] when crossing an increasingly large Bose-glass region in the phase diagram.
7 Conclusion
@@
In summary, we identified the Bose-glass phase experimentally in both weakly and strongly interacting regimes in a two-dimensional square lattice with controllable and reproducible disorder potentials. Using Talbot interferometry across the superfluid-to-Bose-glass transition, we observed a change from long to short-range coherence. Additionally, we found that transitioning in and out of the Bose glass is non-adiabatic on the accessible experimental timescales. A similar behavior was also observed in a quasicrystalline lattice at the superfluid-to-Bose-glass transition [68], which opens opportunities to compare and contrast the behavior of these systems and random disorder on Bravais lattices. It may be possible to cross transitions adiabatically on longer timescales, by reducing trap losses and heating, via ultra-low vacuum pressure [37], and optical lattices with reduced intensity noise. This would enable the exploration of ergodic and non-ergodic phases in a single model, for both weak and strong interactions, through scaling of the adiabatic timescales for traversing transitions or restoring system properties.
@@
Our low-energy studies, combined with Talbot interferometry provide a starting point to further explore disordered systems in and out of equilibrium, and particularly to probe the timescales to establish phase coherence in these systems. This is of relevance to understanding disordered and glass-like quantum states which have been observed in, e.g., cavity polaritons [36], quantum magnets [69] and overdoped cuprates [64]. Future studies with ultracold atoms could focus on the influence of disorder in different lattice geometries [26, 38] and lower dimensions [17, 49], or further explore the stability of many-body localization [18] and the onset of quantum avalanches [34]. Furthermore, controlled disorder is important in studying frustration and stability of the spin-glass phase in the Fermi-Hubbard model [65, 53, 4, 43]. By extending the system sizes is possible to study the similarities and differences between disordered and quasicrystalline systems further, e.g., by realization of the 2D extension of the Aubry-Andre model [61, 21], which could feature mobility edges, critical states and give further insights into the nature of many-body localization and the Bose glass [68].
After completing this work, we became aware of a new preprint studying the disordered Bose-Hubbard model using a superconducting processor [63].
Acknowledgments
Author contributions
LK and CP conducted the experiments and analyzed the data, with support from TH, EH, ALR and SK. CD, PK and AD developed the Quantum Monte-Carlo simulations, while TH led the simulations of the Talbot interferometry. All authors contributed to interpreting the results and writing the manuscript.
Competing interests
There are no competing interests to declare.
Data availability
The data used in this publication is openly available at the University of Strathclyde’s KnowledgeBase [1].
Appendix
@@ Acknowledgments
We use the same quantum-gas microscope setup as detailed in our previous work [19]. Initially, we prepare a Bose-Einstein condensate of $\sim 200$ atoms in a single antinode of an optical lattice (wavelength $\lambda=1064\,\mathrm{nm}$) in the vertical direction, with a depth of $V_{z}=25\,E_{\rm{r}}$. To prepare a disordered system in the weakly interacting regime, we turn on the horizontal lattices, $V_{x}$ and $V_{y}$, to the desired depth, $V_{f}\leq 10\,E_{\rm{r}}$, using an s-shaped ramp within 200 ms and then turn on the disorder potential within 200 ms again using an s-shaped ramp. To prepare a disordered system in the strongly interacting regime, we first turn on the horizontal lattices to reach $V_{x,y}=10\,E_{\rm{r}}$, before adding the disorder while simultaneously increasing $V_{x}$ and $V_{y}$ to their final values, $V_{f}>10\,E_{\rm{r}}$ in a further $200\,\mathrm{ms}$. We detect the atoms by collecting their fluorescence light with a high-NA microscope objective and use the Lucy-Richardson deconvolution method [31] to infer the site occupations.
@@
We create the repulsive disorder potentials using a laser ($\lambda=666\,\mathrm{nm}$) incident on a digital micromirror device (DMD), similar to our previous work [19]. The light from the DMD is overlapped with the optical path used for fluorescence imaging of the atoms and projected through the microscope onto the atoms in the 2D plane. The magnification is such that approximately $18\times 18$ pixels on the DMD correspond to one lattice site. We create a potential at each lattice site (pictured in Fig. 1b), with random amplitude $\Delta_{i}\in[0,\Delta]$, that follows a uniform box distribution. The disorder pattern covers a region with a diameter of 25 lattice sites.
A motorized flip mirror can direct the light pattern from the DMD onto a monitor camera. We employ an iterative intensity feedback algorithm to shape the Gaussian beam incident on the DMD into the target disorder pattern. The light pattern is blurred by the microscope objective when it is projected onto the atoms. To compensate for this, we convolve the monitor camera images with the point-spread function of the objective to effectively optimize the potential as seen by the atoms. Once the DMD pattern has been calculated, it is saved and displayed continuously. The disorder strength is controlled solely by changing the laser power. As detailed in our previous work [19], we measure the phase drift of the optical lattice between experimental realizations by analyzing the position of single atoms in the fluorescence image. This information is then used to shift the position of the DMD pattern for the next measurement, ensuring that the disorder potential remains aligned with the optical lattice. The estimated phase drift from one experimental realization to the next is $0.05\,a_{\rm{\ell}}$.
@@
Experimentally, the thermalization in two dimensions after the one-dimensional Talbot evolution averages over the different chain lengths in our system. We numerically checked that explicitly adding the energy of the one-dimensional chains of different length results in the same behavior as Eq. 6. We have compared results with and without the trapping potential of the vertical lattice and found no significant effect over the timescales probed in the experiment.
Supplemental Materials
@@ 4 Additional data: Talbot interferometry
5 Additional data: Entering the Bose-glass phase